# Quaternion tensor low-rank approximation using a family of non-convex norms

### v3

@@ 1 Introduction
Low-rank matrix approximation (LRMA) is an emerging mathematical tool widely used in various real-world applications, including image Denoising, Inpainting, and Deblurring. LRMA variantes primarily rely on two approaches: matrix factorization and matrix rank. Matrix factorization decomposes a matrix into two or more thinner matrices. A well-known example is Non-negative matrix factorization [20], which decomposes the original Data matrix into two factor matrices with non-negative constraints, whose product closely approximates the original matrix. Numerous extensions of LRMA have been developed, such as robust principal component analysis (RPCA) [5] and matrix completion LRMC [2], among others. This paper focuses on LRMC and RPCA.
The LRMC problem has been extensively researched and shown to be highly beneficial in Completion tasks. The problem in computer vision and graphics is known as image and video Inpainting problem. Typically, this involves reconstructing missing data from available data by assuming the data is low-rank. However, most image and video recovery models are designed for gray scale images, while color images and videos require more complex processing. In the context of processing color videos and images, conventional matrix-based techniques typically overlook. Low-rank tensor comes into play as an extension of the work, allowing color images to be represented as three-dimensional tensors rather than reshaped matrices, and video to four-dimensional tensors. This results in the ability to encode a colored image in a third order tensor, without the need to work only matrices, by converting to gray sclae the image, or reshaping (stacking height and width in single column). Analogously, this goes the same for Hyperspectral image, Video…
A similar approach to handle the high dimensional Data to recover a low rank Data, along with a sparse components, both accurately, is the Tensor Robust Principal Component Analysis (TRPCA) [15, 23], which is an extension of the RPCA, that is also considers as one the first polynomial extension of the Principal Component Analysis. In this model, there is no pre-Set given for known Data, there has been multiple work in this domain.
Recent works make use of the quaternion to represent a colored image, or a video, it has been shown to get a good result in Denoising, histopathological image analysis, color object detection [29, 10] …, as well as in completion and object detection [21]. The quaternion domain allows to encode the red, green and blue channel pixel values on a single entry, this representation can inherent the color structure well, and captures the correlation between the channels, which helps to the success of these methods.
Since the rank minimization is NP-hard, most methods use the the nuclear norm, as it is [4] the tightest convex relaxation of the rank function. However, although the problem is convex and easy to solve, it might not well approximates the rank [7] and gives sub-optimal results. This is due to the fact that, despite the largest singular values can contain more information, each singular value is handled equally.
Several Non-Convex surrogate functions of rank, including the Geman [12] and Laplace norm [9], summarized in Table 1, have been proposed in an attempt to better approximate the rank function. These functions have demonstrated promising results in gray scale image processing. The use of these Non-Convex surrogate functions, has been used in many areas, as clustering [35], Denoising of Hyperspectral images[6], which has widely application as biomedical imaging, and military surveillance. To overcome the aforementioned issues, we consider a series of methods based on the Non-Convex surrogate functions to approximate the low rank quaternion tensor in different proposed models, and show its advantages. The main contribution of the paper are summarized as follows.
@@
  • Propose a model based on Non-Convex surrogate functions for low rank quaternion tensor completion problem, named Low-rank Quaternion Tensor Completion via Non-Convex Tucker Rank (LRQTC-NCTR).
  • Give theoretical analysis of the local convergence of the Completion method.
@@
  • Introduce a second model based on the TT-rank, named Low-rank Quaternion Tensor Completion via Non-Convex TT-rank (LRQTC-NCTTR).
@@
  • Present a third model for Denoising based on Non-Convex surrogate functions, named Tensor Robust Principal Component Analysis via Non-Convex Norms (TRPCA-NC).
  • Demonstrate the advantages of the proposed methods in completing colored videos and in Denoising applications compared to similar methods.
@@
The rest of this paper is structured as follows. Section 2 introduces some preliminaries of quaternion tensors, 3 introduces the the Completion problem, and 4 introduces the second method for the Denoising problem. We propose an extension to the proposed methods in Section 5 before we evaluate the performance of the method in Section 6, and conclude this paper in Section 7.
@@ 2.2 Quaternion matrix and tensor
We denote $\mathbb{H}^{I_{1}\times\ldots\times I_{N}}$ the collection of $I_{1}\times\ldots\times I_{N}$ multi-array (tensors) with quaternion entries. An N-th order quaternion tensor $\dot{\mathcal{Q}}$ is represented as $\dot{\mathcal{Q}}=\mathcal{Q}_{0}+\mathcal{Q}_{1}\mathbf{i}+\mathcal{Q}_{2}\mathbf{j}+\mathcal{Q}_{3}\mathbf{k}$, with $\mathcal{Q}_{i}\in\mathbb{R}^{I_{1}\times\ldots\times I_{N}},i\in[|0,3|]$. Another form that is well used, called the Cayley-Dickson Form, which represents the tensor as sum of two components, i.e, $\dot{\mathcal{Q}}=\mathcal{Q}_{a}+\mathcal{Q}_{b}\mathbf{j}$, with $\mathcal{Q}_{a}=\mathcal{Q}_{0}+\mathcal{Q}_{1}\mathbf{i},$ and $\mathcal{Q}_{b}=\mathcal{Q}_{2}+\mathcal{Q}_{3}\mathbf{i}$.
@@
Given the QSVD $\dot{Q}=\dot{U}S\dot{V}^{H}$, with $S=\operatorname{diag}(\sigma_{i})$. The nuclear norm of $\dot{Q}$ is $\dot{Q}||_{*}:=\sum_{i}\sigma_{i}$, i.e, the sum of its singular values.
@@ 2.3 The QT-product
and $\widehat{(.)}$ is the $(\dot{M}_{i})_{i=3,\ldots,N}$ transformation, i.e,
@@
  • $\dot{\mathcal{Q}}$ is f-diagonal if its frontal slices are diagonal.

### v4

@@ 1 Introduction
Low-rank matrix approximation (LRMA) is a fundamental mathematical tool used in a wide range of real-world applications, including image denoising, inpainting, and deblurring. LRMA techniques generally fall into two broad categories: matrix factorization-based approaches and matrix rank minimization methods. Matrix factorization decomposes a matrix into lower-dimensional factors, improving computational efficiency and reducing noise. A well-known example is Non-negative Matrix Factorization (NMF) [20], which approximates a data matrix using two non-negative factor matrices. Another prominent method is Robust Principal Component Analysis (RPCA) [5], which decomposes a matrix into a low-rank component and a sparse component, enhancing robustness to outliers. Low-Rank Matrix Completion (LRMC) [2] has also been widely studied for recovering missing entries from partially observed matrices under a low-rank assumption.
This paper focuses on LRMC and RPCA, particularly their limitations in handling high-dimensional and color data. The LRMC problem is particularly relevant in computer vision and graphics, where it is commonly referred to as the image and video inpainting problem. The objective is to reconstruct missing or occluded content by exploiting the assumption that the underlying data has a low-rank structure. However, most existing recovery models are primarily designed for grayscale images, while color images and videos require more sophisticated handling due to their higher dimensionality and inter-channel correlations. Traditional matrix-based approaches often ignore essential structural information in color data. For instance, color images are typically represented by three separate RGB channels. Conventional methods often convert these to grayscale or flatten them into matrices, which can result in information loss. A more effective approach is to model color images as third-order tensors and videos as fourth-order tensors. This representation preserves spatial and spectral correlations and avoids unnecessary reshaping operations. Similarly, hyperspectral images, which contain multiple spectral bands, are naturally modeled as higher-order tensors. To further enhance low-rank modeling, Tensor Robust Principal Component Analysis (TRPCA) [lu2016tensor, 15, 23] extends RPCA to third-order tensors. TRPCA decomposes a tensor into a low-rank component and a sparse component, supporting applications such as video inpainting and denoising. Unlike matrix-based RPCA, TRPCA employs the tensor nuclear norm (TNN) within the t-SVD framework, enabling more effective tensor rank minimization. An advantage of TRPCA is that it does not require prior knowledge of the rank or sparsity levels. Convex optimization techniques, such as the Alternating Direction Method of Multipliers (ADMM), are typically used to solve the resulting problem efficiently.
Recent work has shown that quaternion representations are particularly effective for color images and videos. By encoding RGB channels as quaternion entries, quaternion-based methods preserve color structure and inter-channel dependencies. These models have demonstrated success in various vision tasks, including denoising, histopathological image analysis, and color object detection [29, gai2015denoising]. Minimizing the rank function is an NP-hard problem. A common approach is to use the nuclear norm as a convex surrogate [4], but it treats all singular values equally and may fail to approximate the true rank accurately [7]. To overcome this, several non-convex surrogate functions have been introduced, including the Geman function [12] and Laplace penalty [9], which have shown improved results in grayscale image recovery. To address the limitations of existing approaches, this paper introduces a series of quaternion-based low-rank tensor models using non-convex surrogate functions. The key contributions of this work are:
@@
  • A novel model for quaternion tensor completion using non-convex Tucker rank minimization, named Low-Rank Quaternion Tensor Completion via Non-Convex Tucker Rank (LRQTC-NCTR).
  • A theoretical analysis of the local convergence properties of the completion method.
@@
  • A second model for quaternion tensor completion based on TT-rank, named Low-Rank Quaternion Tensor Completion via Non-Convex TT-Rank (LRQTC-NCTTR).
@@
  • A third model for denoising using non-convex surrogate functions, named Tensor Robust Principal Component Analysis via Non-Convex Norms (TRPCA-NC).
  • Comprehensive experiments demonstrating the superiority of the proposed methods for color video completion and denoising applications compared to conventional approaches.
@@
The remainder of this paper is organized as follows. Section 2 introduces some preliminaries of quaternion tensors, Section 3 introduces the Completion problem, and Section 4 introduces the second method for the Denoising problem. We propose an extension to the proposed methods in Section 5 before we evaluate the performance of the method in Section 6, and conclude this paper in Section 7.
@@ 2.2 Quaternion matrix and tensor
We denote $\mathbb{H}^{I_{1}\times\ldots\times I_{N}}$ the collection of $I_{1}\times\ldots\times I_{N}$ multi-array (tensors) with quaternion entries. An $N$-th order quaternion tensor $\dot{\mathcal{Q}}$ is represented as $\dot{\mathcal{Q}}=\mathcal{Q}_{0}+\mathcal{Q}_{1}\mathbf{i}+\mathcal{Q}_{2}\mathbf{j}+\mathcal{Q}_{3}\mathbf{k}$, with $\mathcal{Q}_{i}\in\mathbb{R}^{I_{1}\times\ldots\times I_{N}},i\in[|0,3|]$. Another form that is common to use, called the Cayley-Dickson Form, which represents the tensor as sum of two components, i.e., $\dot{\mathcal{Q}}=\mathcal{Q}_{a}+\mathcal{Q}_{b}\mathbf{j}$, with $\mathcal{Q}_{a}=\mathcal{Q}_{0}+\mathcal{Q}_{1}\mathbf{i},$ and $\mathcal{Q}_{b}=\mathcal{Q}_{2}+\mathcal{Q}_{3}\mathbf{i}$.
@@
Given the QSVD $\dot{Q}=\dot{U}S\dot{V}^{H}$, with $S=\operatorname{diag}(\sigma_{i})$. The nuclear norm of $\dot{Q}$ is $\dot{Q}||_{*}:=\sum_{i}\sigma_{i}$, i.e., the sum of its singular values.
@@ 2.3 The QT-product
and $\widehat{(.)}$ is the $(\dot{M}_{i})_{i=3,\ldots,N}$ transformation, i.e.,
@@
  • $\dot{\mathcal{Q}}$ is f-diagonal if its frontal slices in the transformed domain are diagonal.